BioInformatik (SS 2000)

01/06

Playing GO with an Artificial Neural Net Simulator

Nicolas Goell (nic-goell@magnet.at) Roland Knall (roland.knall@gmx.net) Mark Rehbichler (mark-rehbichler@magnet.at)

Gruppe 2